OpenStudy (anonymous):

given that (5,4) is on the graph of f(x) find the corresponding point for the function f(-1/3x)

OpenStudy (anonymous):

\[f \left(-\frac{ 1 }{ 3 } x\right)\]

OpenStudy (anonymous):

please help I really do not understand this at all

zepdrix (zepdrix):

If (5,4) is on the graph... then the function f(x) evaluated at x=5, will give us an output of 4.

zepdrix (zepdrix):

Meaning, f(5)=4

zepdrix (zepdrix):

Hmmm what can we do with that :d Sec, trying to understand the question.

zepdrix (zepdrix):

@campbell_st @peachpi

OpenStudy (anonymous):

F(-1/3x) The -1/3 will multiply our x values by -3.. Does that make sense?

OpenStudy (anonymous):

I think so, you multiply 5 and 4 by -3?

OpenStudy (anonymous):

so to get (5.4) in this new equation you'd need.. F(-1/3x) x=-15 No your y does not change only how fast or slow it takes for the graph to reach that point

OpenStudy (anonymous):

if it was 3f(-1/3x) then we would multiply our "y" 3 times and our x by -3

OpenStudy (anonymous):

okay now I'm confused

zepdrix (zepdrix):

oh is that what they wanted? Hmm wording was really confusing :U

OpenStudy (anonymous):

All I know is I'm supposed to enter an ordered pair and x is supposed to stay the same

OpenStudy (anonymous):

Suppose, F(x) = 3x F(5) = 3(5) F(5)= 15 (5,15)

OpenStudy (anonymous):

Now we have F(-1/3x) what would be the corresponding point for the function

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

so we know when x=5 then we have 15

OpenStudy (anonymous):

f(-1/3x) = F(5) -1/3x =5 x=-15

OpenStudy (anonymous):

so that's it?

OpenStudy (anonymous):

so on our second equation we know if x is -15 we have f(5) = y

OpenStudy (anonymous):

F(-1/3*-15) F(5) =4

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

its just your x changing not y

OpenStudy (anonymous):

okay so 5 stays the same, what happens to y?

OpenStudy (anonymous):

So we have a point (5,4)

OpenStudy (anonymous):

5,4 is the original points given, I didn't even think it was possible that the answer would literally be in the question

OpenStudy (anonymous):

this can be rewritten as f(5)= 4 F(x) we have this new equation which is the same general shape but it has f(-1/3)

OpenStudy (anonymous):

F(-1/3x)

OpenStudy (anonymous):

(